OpenStudy (anonymous):

Find \(\large \left(\frac{2}{6}\right)^2 \cdot \left(\frac{9}{27}\right)^3\).

ganeshie8 (ganeshie8):

\(\huge \left(\frac{2}{6}\right)^2 \cdot \left(\frac{9}{27}\right)^3\)

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

look at the second fraction, 9 goes in 27 right ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

how many times ?

OpenStudy (anonymous):

3

ganeshie8 (ganeshie8):

\(\huge \left(\frac{2}{6}\right)^2 \cdot \left(\frac{9}{9*3}\right)^3\) \(\huge \left(\frac{2}{6}\right)^2 \cdot \left(\frac{1}{3}\right)^3\)

ganeshie8 (ganeshie8):

Oh look at first fraction, 2 also goes in 6, right ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

\(\huge \left(\frac{2}{6}\right)^2 \cdot \left(\frac{9}{9*3}\right)^3\) \(\huge \left(\frac{1}{3}\right)^2 \cdot \left(\frac{1}{3}\right)^3\)

ganeshie8 (ganeshie8):

that simplifies to : \(\huge \frac{1}{3^2 \bullet 3^3}\)

ganeshie8 (ganeshie8):

further simplifies to : \(\huge \frac{1}{3^{2+3} }\)